OpenStudy (anonymous):

i need to write the vertex form of the following parabola: 1/2(y+4)=(x-7)^2

OpenStudy (anonymous):

put x-7=0 and y+4=0 solve and you get the vertex.

OpenStudy (anonymous):

\[{1\over2}(y+4)=(x-7)^2\]\[y+4=2(x-7)^2\]\[y=2(x-7)^2-4\]

OpenStudy (anonymous):

what you get same thing from my calculation x=7 y=-4 vertex (7,-4)

OpenStudy (anonymous):

yes but why would you do that when the vertex is given in the equation..

OpenStudy (anonymous):

the form it was in:\[{1\over a}(y-k)=(x-h)^2\]where \((h,k)\) is the vertex the form we need is the standard form:\[y=a(x-h)^2+k\]you literally just rearrange it :)

OpenStudy (anonymous):